OpenStudy (anonymous):

Please help!! :) If GM = 13 and GR = 20 what would AR be? Parallelogram in comments..

OpenStudy (anonymous):

Here it is. Idk how to post it in the questing. Srry.

OpenStudy (anonymous):

\(AR=GM\) i guess

OpenStudy (anonymous):

that would make it \(13\)

OpenStudy (anonymous):

Oh duh! It's a parallelogram >.< Sorry. I'm and idiot..

OpenStudy (anonymous):

lol i guess it is a parallelogram, otherwise you can't solve anything

OpenStudy (anonymous):

@satellite73 Yeah true! :p Thx for the help :) Do you know how to solve this by chance? If M∠G = 2x + 11 and m∠A = 5x - 37, calculate M∠G

jimthompson5910 (jim_thompson5910):

If this is a parallelogram, then the opposite angles are congruent (ie equal)

jimthompson5910 (jim_thompson5910):

So M∠G = m∠A 2x+11 = 5x-37 x = ??

OpenStudy (anonymous):

Idk how to solve that >.< I'm reaaaally bad at math, srry.

jimthompson5910 (jim_thompson5910):